A circuit consists of a resistor, capacitor, and inductor connected in series to an AC source. As the source frequency increases, the current in the circuit decreases. Which statement about the circuit is NOT correct as the source frequency increases?

a) The inductive reactance increases.

b) The circuit is said to become more capacitive than inductive.    

c) The total power from the source decreases.

d) The impedance of the circuit increases.

e) The phase angle for the circuit becomes more positive.
Answer #1

a) Inductive reactance X_L = 2pi*f*L , f is frequency. So inductive reactance increases with frequency.

b) Character of circuit depends on Reactance of capacitor and inductor.

Capacitive reactance X_c = 1/2pi*f*C . it is decreasing with frequency increase.

So circuit is being more inductive.

c)Total Power is decreasing because current is derceasing. Power P = (I^2)R, R is resistance(constant)

d) Impedence of circuit is Z = V/I, Voltage V is constant and Current I is decreased so Z is increased

e) Phase angle Phi = tan-1((X_L - X_c)/R). X_L is increasing so Phase angle is becoming more positive.

Only B) is wrong.